OpenStudy (anonymous):

Let f(x) be the function defined below, for the interval -1 < x < 1. f(x) = 7 + x^2 + tan(pi*x/2) (a) find f-1(7) (b) find f(f-1(4)) Note: f-1 stands for inverse.

OpenStudy (anonymous):

How would I go about with doing this?

OpenStudy (anonymous):

We have to realize or work with the definition of inverse in mind. Keeping that in mind, realize that if the f^-1(7) then f(y)=7. Thus we have to find the value that makes the original equation 7. Namely we have to solve: 7+x^2+tan(pi*x/2)=7

OpenStudy (anonymous):

That's exactly what I came up with in the beginning. Only right after that, I realized that I don't know how to proceed from there.

OpenStudy (anonymous):

subtract 7 from both sides

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then, its acutally quite comical how we can , just using out eyes realize when x^2 and the tan fucntion are 0

OpenStudy (anonymous):

im sorry but im not seeing it unfortunately.

OpenStudy (anonymous):

come on man, just look at it: x^2+tan(pi*x/2)=0 When is x^2 0? When is tan(pix/2) 0?

OpenStudy (anonymous):

at x = 0, at x^2 = 0. at x = 0, tan(pi*x/2) also equals 0.

OpenStudy (anonymous):

so the f^-1(7)=0

OpenStudy (anonymous):

So how would part b work?

OpenStudy (anonymous):

simply: f(g(x)=x Where g(x) is the inverse

OpenStudy (anonymous):

you just got schooled

OpenStudy (anonymous):

We don't have the g(x)

OpenStudy (anonymous):

do i have to spell it out, i guess so: f(f^-1(4))=4

OpenStudy (anonymous):

Aw man you shouldn't have. Ok thanks anyways.